OpenStudy (anonymous):

Put the following equation in slope intercept form. 10x - 5y = -20 Is the correct answer y = 2x + 4? True False

sam (.sam.):

yes

sam (.sam.):

10x-5y=-20 -5y=-10x-20 y=2x+4